OpenStudy (anonymous):

60 percent of what number is 15 60 is 15% of what number

hero (hero):

Use this formula to figure it out: \[\frac{\text{part}}{\text{whole}} = \frac{\%}{100}\]

OpenStudy (anonymous):

well for the second part you would multiply 60 by 15% to get the number

hero (hero):

For the first one: \[\frac{15}{x} = \frac{60}{100}\]

hero (hero):

If you simplify the right fraction you get: \[\frac{15}{x} = \frac{3}{5}\] Then cross multiply to solve for x.

OpenStudy (anonymous):

then for the first one you would do 15 devided by 60 to get an number and then you multiply that number by 100 to get the answer

OpenStudy (anonymous):

*divided